### v2

# A Novel Approach to Reduce Derivative Costs in Variational Quantum Algorithms

Abstract
We present a detailed numerical study of an alternative approach, named Quantum Non-Demolition Measurement (QNDM) Solinas et al. (2023), to efficiently estimate the gradients or the Hessians of a quantum observable. This is a key step and a resource-demanding task when we want to minimize the cost function associated with a quantum observable. The statistical study of the error leads to a further performance improvement with respect to the original proposal in Ref. Solinas et al. (2023). In our detailed analysis, we account for all the resources needed to implement the QNDM approach with a fixed accuracy and compare them to the current state-of-the-art method Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We find that the QNDM approach is more efficient, i.e. it needs fewer resources, in evaluating the derivatives of a cost function. These advantages are already clear in small dimensional systems and are likely to increase for practical implementations and more realistic situations. A significant outcome of our study is the implementation of the QNDM method in Python, provided in the supplementary material S.Caletti and G.Minuto (2024). Given that most Variational Quantum Algorithms can be formulated within this framework, our results can have significant implications in quantum optimization algorithms and make the QNDM approach a valuable alternative to implement Variational Quantum Algorithms on near-term quantum computers.
I Introduction
Complex optimization problems, such as drug, molecular, and material design, are some of the research fields in which quantum computers are likely to have an impact in the mid-term timescale Gilyén et al. ; Moll et al. (2018); Barkoutsos et al. (2021). These problems share a common structure: given a cost function that represents our physical quantity of interest, we would like to find its minimum. The scheme to approach these problems in a quantum computer is a hybrid quantum-classical one, where Variational Quantum Algorithms (VQA) McClean et al. (2016) can be naturally implemented. The picture is the following: quantum computers are used to calculate the gradient of the cost function evaluated at a certain point in the parameter space. This can be done with different techniques but is usually done by evaluating the cost function in two points that are close in the parameter space and exploiting the parameter-shift rule to compute the derivative Mari et al. (2021). Then, this information is fed into a classical computer, which calculates how to adjust the parameters of the quantum circuit in order to reduce the cost function. By iterating this procedure, as in variational problems in physics, the quantum circuit should converge toward the configuration that generates one of the states minimizing the cost function (either a local or a global minimum).
The classical optimization procedures are well-known and established Boyd and Vandenberghe (2004); Adby (2013) and they usually exploit information on the derivatives (gradient) of the cost function to update the parameters. At the quantum level, the way to extract the desired information from a quantum system is still debated and open Banchi and Crooks (2021); Wierichs et al. (2022). Indeed, the derivative cannot be directly associated with a Hermitian operator Solinas and Gasparinetti (2015, 2016) and, therefore, there is no unique way to measure it. Still, this is a crucial step since it usually causes a bottleneck in computation with quantum computers. The most straightforward and used approach to obtain the gradient of the cost function is to run a quantum circuit and measure a quantum observable at a certain point in the parameter space. Then, we repeat the procedure changing the circuit parameters by a little along each direction, separately Mari et al. (2021); Schuld et al. (2019); Cerezo et al. (2021). We call this method, the Direct Measurements (DM) approach and, for every point in the parameter space, it requires two observable measurements to determine each component of the gradient.
In this article, we analyze a novel method to compute the derivatives of the cost function implemented on the quantum circuit. This is called the Quantum Non-Demolition Measurement (QNDM) approach and it was first theoretically described in Ref. Solinas et al. (2023). The main idea is to extract the information about the derivative with a single measurement of a quantum observable. This is obtained by coupling a quantum detector to the original quantum system so that the information about the value of the derivative of the cost function is stored in the phase of the detector. Finally, the phase is measured to extract this information Solinas et al. (2023).
In the following, we present a detailed comparison between the two methods. We consider the evaluation of the cost function at a fixed accuracy, and we extract information about the repetitions of the circuit evaluation required. We do that by estimating the statistical error (Mean Squared Error) in the two approaches, and we calculate the total resource cost in both cases. This theoretical analysis is followed by full computational simulations, in which we compare the total resources needed in the two approaches to estimate the derivative of the cost function directly from their implementations.
We find that, already for cases with limited complexity, the QNDM approach has a substantial advantage over the DM ones. The resource-saving will increase for more practical and interesting problems of intermediate dimension.
The paper is structured as follows. In Sec. II, we recall the basic ideas of VQA and the two approaches to derivative evaluation. Section III is devoted to the analysis of the statistical errors and the resource cost. Section IV.2 presents the numerical comparison, while Sec.VI contains the conclusions. The codes to calculate derivatives with QNDM used within this article are accessible via GitHub S.Caletti and G.Minuto (2024). In Appendix B we show how to install the QNDM package.
II Theoretical Discussion
II.1 General Settings
We briefly recall the standard implementation of Variational quantum algorithms in a quantum computer (for more details see Refs. McClean et al. (2018); Mari et al. (2021); Solinas et al. (2023)). We consider a system of $n$ qubits initialized in the state $|\psi_{0}\rangle=|0\rangle_{1}\otimes|0\rangle_{2}\otimes\ldots\otimes|0\rangle_{n}\equiv|00...0\rangle$. The generic unitary transformation acting on them is denoted $U({\vec{\theta}})$, where $\vec{\theta}$ is a vector in the parameter space. The quantum state obtained by applying the transformation $U({\vec{\theta}})$ to the initial state of the circuit is given by $|\psi({\vec{\theta}})\rangle=U({\vec{\theta}})|\psi_{0}\rangle$. The $U({\vec{\theta}})$ transformation can be implemented as a sequence of parameterized unitary transformations $U_{j}({\theta_{j}})$ and arbitrary transformations (independents of $\vec{\theta}$) $V_{j}$, $j\in[1,m]$, such that
$U({\vec{\theta}})=V_{m}U_{m}({\theta_{m}})....V_{2}U_{2}({\theta_{2}})V_{1}U_{1}({\theta_{1}})\,.$ (1)
We consider $U_{j}({\theta_{j}})=\exp{\{-i\theta_{j}\hat{H}_{j}\}}$ and $\hat{H}_{j}$ is the corresponding generator of the transformation. The quantum circuit to generate this transformation is represented in Fig. 1 a). Assuming that $\hat{H}_{j}^{2}=\mathds{1}$, we have that
$U_{j}(\theta_{j})=e^{-i\hat{H}_{j}\theta_{j}/2}=\cos\frac{\theta_{j}}{2}\mathds{1}-i\sin\frac{\theta_{j}}{2}\hat{H}_{j}.$ (2)
This accounts for a multitude of relevant cases such as when $H_{j}$ is a tensor product of any multi-qubits Pauli matrices Mari et al. (2021). However, to further simplify the discussion and the simulation, in the following, we restrict our attention to the case (presented in Fig. 1) in which
$U_{j}(\vec{\theta}_{j})=R^{j}_{1}(\theta^{j}_{1})\dots R^{j}_{n}(\theta^{j}_{n})\,,$ (3)
where each operator $R^{j}_{i}(\theta^{j}_{i})$ is a parameterized single qubit gate (see Eq. (2)) and $\hat{H}_{j}$ is a single Pauli matrix. Likewise, for the $V_{j}$ part of the transformation, which is independent of the parameters, we select the following specific structure
$V_{j}=C_{1}NOT_{2}\dots C_{n-1}NOT_{n}\,,$ (4)
where $C_{i-1}NOT_{i}$ is a two-qubit operator with the control on the $i-1$-th qubit and the action on the $i$-th qubit Nielsen and Chuang (2010). The sequence of $U_{j}(\vec{\theta}_{j})$ and $V_{j}$ is usually called layer and, as shown in Fig. 1, is a combination a parametrized and entangling unitary transformations.
Figure 1: a) Structure for the $j$-th layer as a sequence of a parameterized layer $U_{j}(\vec{\theta}_{j})$ and an unparameterized layer $V_{j}$ for a $n$ qubit system. b) Example of a layer in terms of single-qubit gates depending on the $\theta_{i}^{j}$ parameters and two qubits gates $C_{i-1}NOT_{i}$.
In optimization problems, the goal is to find the quantum state $\mathinner{|{\psi(\vec{\theta})}\rangle}$ which minimizes the expectation value of a quantum observable $\hat{M}$. To accomplish this task, we define a cost function $f$ that we want to minimize with respect to the parameters ${\vec{\theta}}$
$f({\vec{\theta}})={\mathinner{\langle{\psi({\vec{\theta}})}|{\hat{M}}|{\psi({\vec{\theta}})}\rangle}}={\mathinner{\langle{0}|{U^{\dagger}({\vec{\theta}})\hat{M}U({\vec{\theta}})}|{0}\rangle}}\,.$ (5)
In many applications McArdle et al. (2020); Tilly et al. (2022), the observable $\hat{M}$ can be written as a weighted sum of Pauli strings $\hat{P}_{i}=\prod_{j}\hat{A}_{i}^{j}$, where $\hat{A}_{j}\in[X_{j},Y_{j},Z_{j},I_{j}]$ is one of the usual Pauli operators acting on the $j$-th qubit. With this notation, we have that
$\hat{M}=\sum_{i=1}^{J}h_{i}\hat{P}_{i}$ (6)
where $J$ is the total number of Pauli string composing $\hat{M}$.
We adopt the so-called hybrid quantum-classical optimization procedure described in the introduction McClean et al. (2016); Cerezo et al. (2021). The minimization algorithm consists of two steps that must be iterated until the minimum of $f({\vec{\theta}})$ is reached. First, with the help of a quantum computer, we evaluate the gradient of $f({\vec{\theta}})$ in a given point of the parameter space $\vec{\theta}$. Then, the information about the gradient is fed into a classical computer which, with standard optimization algorithms Kingma and Ba (2017); Kübler et al. (2020); Sweke et al. (2020), evaluates in which direction of the $\vec{\theta}$ space we move to reach the minimum of $f$. In this paper, we focus on the quantum part of the algorithm, that is the calculation of the gradient of $f({\vec{\theta}})$ done with a quantum computer and we do not discuss the classical one.
To measure the derivative of $f({\vec{\theta}})$ along the $j$-th direction in the parameter space $\vec{\theta}$, we measure $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$ where $\hat{e}_{j}$ is the unit vector along the $\theta_{j}$ direction and $s$ is a shift parameter. Then, we calculate the quantity Mari et al. (2021)
$g_{j}=\frac{\partial f({\vec{\theta}})}{\partial\theta_{j}}=\frac{f(\vec{\theta}+s\hat{e}_{j})-f(\vec{\theta}-s\hat{e}_{j})}{2\sin s}\,.$ (7)
With $s=\pi/2$, we obtain the parameter shift rule described in Refs. Li et al. (2017); Mitarai et al. (2018); Schuld et al. (2019); Mitarai and Fujii (2019). Notice that $g_{j}$ is exactly the derivative of $f({\vec{\theta}})$ with no approximation and repeating the procedure for all the directions $\theta_{j}$ we obtain the gradient of $f({\vec{\theta}})$, which is then fed into the classical optimization algorithm.
II.2 Direct Measurement
The standard approach to measure the value of the derivatives of the cost function (7) is the Direct Measurement (DM) method Schuld et al. (2019). This method consists of measuring separately the average values of the operator $\hat{M}$ in the two points, i.e., $\vec{\theta}\pm s\hat{e}_{j}$, and then calculating the derivative as in Eq. (7). To calculate the values of $f(\vec{\theta}+s\hat{e}_{j})$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s\hat{e}_{j})$ transformation, and then we perform a projective measure for each Pauli string $\hat{P}_{i}$ that appears in the definition (6) of the observable $\hat{M}$ McArdle et al. (2020); Mari et al. (2021). For every Pauli string $\hat{P}_{i}$ in $\hat{M}$, we have to iterate these steps until the statistical accuracy is reached. The same procedure is then implemented for $f({\vec{\theta}-s\hat{e}_{j}})$ and then, the derivative can be calculated as in Eq. (7).
Following this procedure, a single derivative $g_{j}$ in the DM approach can be written as
$g_{j}=\sum_{i}\frac{h_{i}}{2\sin s}\Big{(}{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\rho_{s}^{0}U(\vec{\theta}+s\hat{e}_{j})\big{]}-{\rm Tr}_{S}\big{[}\hat{P}_{i}U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\rho_{s}^{0}U(\vec{\theta}-s\hat{e}_{j})\big{]}\Big{)}\,,$ (8)
where $\rho_{s}^{0}=|\psi_{0}\rangle\langle\psi_{0}|$, $U(\vec{\theta})$ is the operator in Eq. (1) and ${\rm Tr}_{S}$ denotes the trace over all the qubits. An implementation in terms of quantum circuits is shown in Fig. 2.
Figure 2: Pictorial representation of the quantum circuit for the implementation of the DM protocol. The $U(\vec{\theta}\pm s\hat{e}_{j})$ transformations are the ones in Eq. (1). The sums represent the fact that the circuits followed by the measure must be run for different shifts ±$\pm s$ and for every Pauli string ($i=1,...,J$) in the observable $\hat{M}$ as in Eq. (6).
II.3 Quantum Non-Demolition Measurement
In this section, we briefly recall the main characteristics of the QNDM method Solinas et al. (2023). The key idea is to store the information about the derivative of the cost function in the phase of an ancillary qubit named the detector. More specifically, by coupling the original system and the detector twice, it is possible to store the information about both $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$ in the phase of the detector. Then, the detector is measured to extract information about the derivative Solinas et al. (2023).
The advantage of this approach lies in the fact that the couplings between the system and the detector are sequential and the measure on the detector is performed only at the end of the procedure. Therefore, even if the quantum circuit must be iterated to obtain the desired statistical accuracy, we extract the derivative information by performing just a single measurement on the detector qubit instead of the two needed with the DM approach. The details of the QNDM algorithm can be found in Solinas et al. (2023). Here, we summarize only the main steps.
Suppose we are interested in the quantum observable $\hat{M}$ in Eq. (6). We add an ancillary qubit and implement the system-detector coupling through the operator $U_{\pm}=\exp\{\pm i\lambda Z_{a}\otimes\hat{M}\}$ where $Z_{a}$ is a Pauli operator acting on the ancillary detector and $\lambda$ is the system-detector coupling constant. We notice that QNDM clearly offers an elegant way to implement the operator $U_{\pm}$, as the product of single Pauli string operators $\prod_{i}^{J}\exp\{\pm i\lambda Z_{a}\otimes h_{i}\hat{P_{i}}\}$ Solinas et al. (2023). Considering the described quantum circuit, this corresponds to the following transformation (in the full system and detector Hilbert space)
$U_{tot}=e^{i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}+s\hat{e}_{j})U^{\dagger}(\vec{\theta}-s\hat{e}_{j})e^{-i\lambda Z_{a}\otimes\hat{M}}U(\vec{\theta}-s\hat{e}_{j})\,.$ (9)
We consider as initial state for the system plus the detector $\mathinner{|{\psi_{0}}\rangle}(\mathinner{|{0}\rangle}_{D}+\mathinner{|{1}\rangle}_{D})/\sqrt{2}$. After implementing the transformation (9), we measure the accumulated phase $\exp\{i\phi(\lambda)\}$ on the detector degrees of freedom. It can be shown Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022) that this is a quasi-characteristic function $\mathcal{G}_{\lambda}$
$\mathcal{G}_{\lambda}\equiv e^{i\phi(\lambda)}=\frac{\tensor{[}_{D}]{{\mathinner{\langle{0}|{\rho_{D}^{f}}|{1}\rangle}}}{{}_{D}}}{\tensor{[}_{D}]{{\mathinner{\langle{0}|{\rho_{D}^{0}}|{1}\rangle}}}{{}_{D}}}\,,$ (10)
where $\mathinner{|{0}\rangle}_{D}$ and $\mathinner{|{1}\rangle}_{D}$ are eigenstates of the detector operator $Z_{a}$.
Figure 3: Quantum circuit of the implementation of the QNDM protocol. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U({\vec{\theta}}-s\hat{e}_{j})$, $U_{2}=U^{\dagger}({\vec{\theta}}-s\hat{e}_{j})U({\vec{\theta}}+s\hat{e}_{j})$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
Taking the derivatives of $\mathcal{G}_{\lambda}$ with respect to $\lambda$ and evaluating them in $\lambda=0$, we have access to the information of the different moments of the distribution describing the variation of the cost function Solinas et al. (2023); Solinas and Gasparinetti (2015, 2016); Solinas et al. (2021, 2022). Here, we are interested only in the first moment which directly gives us $g_{j}$. Following Ref. Solinas et al. (2023), it can be shown that the first derivative of $\mathcal{G}_{\lambda}$ reads
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}$ $=$ $2{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\hat{M}U(\vec{\theta}+s\hat{e}_{j})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\hat{M}U(\vec{\theta}-s\hat{e}_{j})\rho_{S}^{0}]$ (11)
$=$ $2\sum_{i}h_{i}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s\hat{e}_{j})\hat{P_{i}}U(\vec{\theta}+s\hat{e}_{j})\rho_{S}^{0}-U^{\dagger}(\vec{\theta}-s\hat{e}_{j})\hat{P_{i}}U(\vec{\theta}-s\hat{e}_{j})\rho_{S}^{0}]\,,$
where ${\rm Tr}_{S}$ denotes the trace over the system degrees of freedom. By direct comparison, Eq. (11) is proportional to Eq. (8),
$-i\partial_{\lambda}\mathcal{G}_{\lambda}\Big{|}_{\lambda=0}=2\sin(s)g_{j}.$ (12)
Thus, both the QNDM and the DM approach can be used to compute the derivative of $f(\vec{\theta})$.
In practice, it is convenient to approximate the derivative of the quasi-characteristic function for small $\lambda$, i.e. in the linear regime. In this approximation, the derivative is given by Solinas et al. (2023)
$-i\partial_{\lambda}\mathcal{G}_{\lambda}|_{\lambda=0}=\partial_{\lambda}\phi(\lambda)|_{\lambda=0}\approx\frac{\phi(\lambda)-\phi(0)}{\lambda}=\frac{\phi(\lambda)}{\lambda}$ (13)
where in the last steps we have used the fact that, for symmetry reason, $\phi(0)=0$ (see, for example, Ref. Solinas et al. (2023)).
Then, the phase accumulated by the detector during the interaction with the system might be measured with interferometric techniques. The other logical operations we need are the Hadamard gate $H$ and a phase gate $S$ defined by Nielsen and Chuang (2010)
$H$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (14)
$S$ $=\begin{pmatrix}1&0\\ 0&e^{i\frac{\pi}{2}}\end{pmatrix}=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$ (15)
In terms of circuit implementation, the interferometric measure is built by applying a Hadamard gate, a phase gate, and another Hadamard to the detector and to measuring the populations of the two possible outcomes $P_{0,1}$ (see Fig. 3).
It can be shown (see Ref. Solinas et al. (2023)) that, for small $\lambda$, $\phi(\lambda)=-\arcsin(2P_{0}-1)$. Furthermore, using Eq. (13) the derivative of the cost function $f(\vec{\theta})$ reads
$g_{j}\sim-\frac{\arcsin(2P_{0}-1)}{2\lambda\sin s}\,,$ (16)
where we have added the term $2\sin s$ to the denominator to directly get the value of the derivatives according to the parameter shift rule.
III Mean Squared Error and cost analysis
Understanding how statistical errors affect the estimate of the derivatives and gradient plays a pivotal role in judging the performance and reliability of the different methods presented above. As a measure of the quality of the estimation of the cost function derivative, we consider Mean Squared Error ($MSE$), i.e., the average of the difference between the estimated values $\hat{g}_{j}$ and the exact value $g_{j}$, squared. We introduce the bias and variance as Cox and Hinkley (1974)
$\text{Bias}(\hat{g}_{j})=\mathbb{E}(\hat{g}_{j})-g_{j}\,,$ (17)
$\text{Var}(\hat{g}_{j})=\mathbb{E}[(\hat{g}_{j}-\mathbb{E}(\hat{g}_{j}))^{2}]\,,$ (18)
where $\mathbb{E}(\cdot)$ is the expectation over the statistical distribution of the measurement outcomes. The bias represents a systematic distortion of a statistical result, and it remains present even in the limit of an infinite number of shots $N$. On the other hand, variance quantifies the degree of dispersion, indicating how far a collection of values deviates from its mean value. Unlike the bias, the variance depends on the shot number $N$ Mari et al. (2021). In terms of these quantities, we define the mean square error (MSE) as
$\text{MSE}(\hat{g}_{j})=\text{Bias}(\hat{g}_{j})^{2}+\text{Var}(\hat{g}_{j})\,.$ (19)
We start our discussion by computing the bias and variance for the QNDM method, following the methodology outlined in Ref. Mari et al. (2021). They read
$\text{Bias}_{QNDM}(\hat{g}_{j})=-i\left(\frac{\phi(\lambda)-\phi(0)}{\lambda}-\partial_{\lambda}\mathcal{G}_{\lambda}\right)=\frac{\lambda\partial^{2}_{\lambda}\mathcal{G}_{\lambda}}{2}+\mathcal{O}(\lambda^{2})\,,$ (20)
$\text{Var}_{QNDM}(\hat{g}_{j})=\sigma_{QNDM}^{2}\,.$ (21)
We observe that the QNDM approach, as expected, exhibits bias different from zero arising from the linear approximation employed in Eq. (13) to obtain the derivative of the quasi-characteristic function $\mathcal{G}_{\lambda}$. Therefore, its value is directly proportional to the second derivative of the function $\mathcal{G}_{\lambda}$ multiplied by $\lambda$. To compute the variance $\sigma^{2}_{QNDM}$, we must incorporate the statistical error $\sigma^{2}_{P_{0}}$ associated with the detector population $P_{0}$ that we measure. For $N_{QNDM}$ single qubit measurements of the detector, we obtain $\sigma^{2}_{P_{0}}=\sigma^{2}_{D}/N_{QNDM}$ where $\sigma_{D}$ is the single shot variance obtained measuring the detector. Following the standard error propagation formulas Taylor (1997) for Eq. (16), we obtain $\sigma^{2}_{QNDM}$ from the variance for the detector population, which is given by
$\sigma^{2}_{QNDM}=\frac{\sigma^{2}_{P_{0}}}{4\sin^{2}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}=\frac{\sigma^{2}_{D}}{4N_{QNDM}\sin^{2}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}.$ (22)
Thus, the corresponding MSE is
$\text{MSE}_{QNDM}(\hat{g}_{j})=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma^{2}_{D}}{4N_{QNDM}\sin^{2}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}.$ (23)
Notably, in MSEQNDM, the variance term is directly proportional to $\lambda^{-2}$ while the bias term is proportional to $\lambda^{2}$. Thus, the optimal reduction of the MSE hinges on selecting an appropriate value for $\lambda$. We determined that an optimal balance is achieved for $\lambda=0.1$. With this specific choice of $\lambda$, in the shot regime commonly employed in optimization, $N\in[10^{3}-10^{4}]$ Kandala et al. (2017); Schuld and Killoran (2019), $\sigma^{2}_{QNDM}$ dominates over the bias squared term so that $\text{MSE}_{QNDM}(\hat{g}_{j})\approx\text{Var}_{QNDM}(\hat{g}_{j})=\sigma^{2}_{QNDM}$. Consequently, in the computational section IV, we exclusively consider the contribution from the variance to calculate the MSEQNDM error.
In the DM approach, we need to measure the system qubits twice to obtain $f(\vec{\theta}+s\hat{e}_{j})$ and $f(\vec{\theta}-s\hat{e}_{j})$. Proceeding as above, we have
$\text{Bias}_{DM}(\hat{g}_{j})=0$ (24)
$\text{Var}_{DM}(\hat{g}_{j})=\frac{\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{j})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{j})}{4\sin^{2}s}.$ (25)
The bias contribution for DM vanishes because the parameter shift rule calculates the exact value of derivative Mari et al. (2021). The variance, instead, is composed by the terms $\sigma_{DM}^{2}(\vec{\theta}\pm s\hat{e}_{j})$. Intuitively, they correspond to the evaluation of $f(\vec{\theta}\pm s\hat{e}_{j})$. Following Ref. Mari et al. (2021), we assume that the variance of the measured observable depends weakly on the parameter shift such that $\sigma_{DM}^{2}(\vec{\theta}+s\hat{e}_{j})+\sigma_{DM}^{2}(\vec{\theta}-s\hat{e}_{j})\sim 2\sigma_{DM}^{2}$ for any value of $s$. As we have seen in sec II.2, to obtain the derivatives, we have to measure each Pauli string individually in the parameter space point $\vec{\theta}\pm s\hat{e}_{j}$ (see Eq. (8) and Mari et al. (2021); Cerezo et al. (2021)). Then, the statistical error formula for the estimation of $f(\vec{\theta})$ is given by
$\sigma^{2}_{DM}=\sum_{i=1}^{J}\frac{h^{2}_{i}\sigma^{2}_{s}}{N_{DM}}$ (26)
where $h_{i}$ are the parameters in $\hat{M}$ in Eq. (6) and $\sigma_{s}$ is the single shot variance obtained measuring the $n$ qubits Mari et al. (2021). In Eq. (26), as in Ref. McClean et al. (2018); Mari et al. (2021), we assume that $\sigma_{s}$ is equal for each Pauli string measurement. Following these observations, we conclude that the MSE of DM approach is only due to the statistical noise and it is equal to
$\text{MSE}_{DM}(\hat{g}_{j})=\sum_{i=1}^{J}\frac{h^{2}_{i}\sigma_{s}^{2}}{2N_{DM}\sin^{2}s}.$ (27)
An important remark must be made about the results obtained in Eq. (23) and (27). The mean square error MSEQNDM does not depend on the number of the Pauli string in $\hat{M}$ or their weights, while MSEDM does. The reason for this is tracked back to the different ways to extract information about the cost function $f$ and its derivatives. In the QNDM protocol all the information about $\hat{M}$ (and the Pauli strings) are stored in series through the system-detector coupling. On the contrary, in the DM protocol, the information about the Pauli strings is retrieved one at a time by running the circuits and performing the measurements. We can estimate exactly the number of shots to get the same precision in MSEDM and MSEQNDM by using Eqs. (23) and (27). We obtain
$\frac{\text{MSE}_{DM}(\hat{g}_{j})}{\text{MSE}_{QNDM}(\hat{g}_{j})}=1\longrightarrow N_{DM}=N_{QNDM}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i},$ (28)
where the $\sum_{i}h^{2}_{i}$ grows proportionally to the number of Pauli strings.
III.1 Cost analysis and comparison between DM and QNDM approaches
To make a comparison between the two approaches, we use the cost $\mathcal{C}\left(g_{j}^{i}\right)$ ($i=QNDM,\,DM$) that gives the scaling in terms of the number of gates required to compute a derivative in the cost function for the two algorithms represented in Figs. 2 and 3. In addition, we use the resource ratio $\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)$.
The resource costs are functions of the number of Pauli strings $J$, the number of qubits $n$, and the number of logical operators $k$ used in the circuits to calculate a single derivative (where in $k$ we include both single-qubit and two-qubit gates). As in the previous section, $N_{DM}$ and $N_{QNDM}$ represent the number of shots or repetitions for DM and QNDM, respectively. The values extracted from Refs. Solinas et al. (2023); Cerezo et al. (2021); Mari et al. (2021) are detailed in Tab. 1.
It is important to note that we assign two different values for the required number of shots because, as discussed in Sec. III, the MSE in the case of DM depends linearly on the number of Pauli strings $J$. Therefore, to obtain the same MSE error in the derivative estimation for QNDM and DM, the number of shots for DM is $N_{DM}\propto N_{QNDM}J$ [see Eq. (28)].
Method Shots Costs
DM $N_{DM}$ $\mathcal{C}\left(g_{j}^{DM}\right)=2N_{DM}J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{j}^{QNDM}\right)=N_{QNDM}(3k+8Jn)$
Table 1: Resource cost function to compute first-order derivatives with the DM and the QNDM approaches.
Given these premises, we are interested in two regimes. First, when $k\gg nJ$, i.e., the number of logical gates to implement $U(\vec{\theta})$ exceeds the number of Pauli strings $J$ (time the number of qubits). From Tab. 1 and Ref. Solinas et al. (2023), we have that the resource ratio scales as
$\mathcal{C}\left(\frac{g_{j}^{DM}}{g_{j}^{QNDM}}\right)=\frac{2J}{3}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i=1}^{J}h^{2}_{i}=\mathcal{O}(J^{2}).$ (29)
This behavior shows that the cost to calculate a derivative with DM is $J^{2}$ higher than QNDM so that, in this regime, we have quadratic speed up with the number of Pauli strings. This arises from two components: firstly, the inherent linear advantage that QNDM  holds over DM  Solinas et al. (2023), and secondly, the dependency of the MSEDM on the number of $J$, which is absent in MSEQNDM. This regime is particularly relevant for quantum chemistry simulations. For instance, in simulations of moderately complex molecules Wecker et al. (2015), we have $k\approx 10^{9}-10^{10}$ operations to implement $U(\vec{\theta})$, $n\approx 10^{2}-10^{3}$ (the expected dimension of quantum computers in the next years), and $J>10^{3}$ as the number of Pauli strings in the Hamiltonian. With these numbers, $k\gg nJ$, resulting in a reduction of roughly $J^{2}$ resources when using the QNDM approach. This advantage is likely to increase for larger and more complex molecules.
This regime holds also significant potential for simulations in quantum machine learning since the expressivity, i.e., the ability of a quantum circuit to generate (pure) states that are well representative of the Hilbert space Sim et al. (2019) of a quantum circuit is directly linked to the number of parametrized rotation in $U(\vec{\theta})$, i.e., $k$.
The opposite regime is reached when $k\ll nJ$, i.e., when the $\hat{M}$ operator is composed of many Pauli strings (thus, its averages might be difficult to estimate or measure) and the logical space to be searched is reduced (thus, we need unitary transformation with a limited number of logical gates). In this case, the ratio between the resources employed reads
$\mathcal{C}\left(\frac{g_{j}^{DM}}{g_{j}^{QNDM}}\right)=\frac{k}{4n}\frac{2\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i}=\mathcal{O}(Jk).$ (30)
Whereas the $k$ factor comes from the linear cost advantage of QNDM with respect to DM for this regime (see Tab. 2), the $J$ remains as a consequence of the different dependency of the MSEDM and MSEQNDM. This scenario is promising for various applications in VQAs as pointed out in Flynn et al. (2022).
IV Computational analysis
IV.1 Errors Analysis
In this section, we present a numerical comparison between errors estimated using the two methods. For the numerical analysis, we have used the simulator provided by IBM Quantum known as Aer IBM Quantum (2024). As noticed at the end of Sec. III, we expect an advantage from the QNDM approach (with respect to the DM approach).
The value of the derivatives $g_{j}$ depends on many parameters:
  1. (i)
    the logical gates needed to implement $U(\vec{\theta})$, e.g., see Eq. (9);
  2. (ii)
    their total number $k$;
  3. (iii)
    the direction $j$ along which is calculated the derivative;
  4. (iv)
    the Pauli strings $\hat{P}_{j}$ that appear in $\hat{M}$;
  5. (v)
    their number $J$, as pointed out in Eq. (6).
For this reason, to ensure statistical significance, we need to average over these parameters. In the following, we use the term “realization” to denote a random choice of all the relevant parameters and possibilities described in the points from (i) to (v).
Figure 4: The main plot represents the value of the derivative over the $j$-th direction $\mu(g_{j}^{i})$ [$i=$DM (red markers) and $i=$QNDM (blue markers)], averaged over a collection of different realizations, and shown as a function of the number of Pauli strings. In the inset, we show the ratio between the $MSE_{DM}/MSE_{QNDM}$. The simulated quantum system has $n=10$ qubits and the number of shots is fixed at $N=10^{4}$. Each point is averaged for $L=100$ realizations implemented as discussed in the main text.
In Fig. 4 we plot the averaged derivative $\mu(g_{j}^{i})$ over a randomly chosen direction $j$, and their errors as a function of the number of Pauli strings $J$. The DM  and QNDM  results correspond to the red and blue markers, respectively. In this case, the average of $\mu(g_{j}^{i})$ is performed over a collection of different realizations averaged over the above parameter (i) to (iv). For each point, we randomly selected $L=100$ realizations. Also, the errors associated with the points are the result of the average of the errors of the corresponding realizations.
Notice that, while the values obtained with the two protocols are compatible within the error bars, the uncertainty in the DM approach increases with $J$ while the QNDM does not. This means that, with the DM approach we must increase the number of shots to reach the same precision obtained with the QNDM one. This is further exemplified by the inset in Fig. 4 where we plot the ratio of the MSEs. As can be seen, the MSEDM is a factor of $10$ larger than MSEQNDM already for $J=100$, and the difference grows linearly with $J$.
IV.2 Costs Comparison
For practical situations, it is of paramount importance to quantitatively estimate any advantage of the QNDM algorithm over DMand in which conditions this is reached. From the discussion in Sec. III.1, we expect the QNDM approach to perform better, i.e., it needs fewer resources, in some specific but interesting regimes. Aiming in this direction, we present a full numerical analysis of the resources needed to calculate the derivative of the cost function.
To have a meaningful comparison, we have fixed the precision, i.e., the MSE required for the derivative evaluation for both approaches. Consequently, for the DM approach, we must increase the number of shots, as we already showed in Eq. (28) and Fig. 4.
Figure 5: Regime $k\gg nJ$. $a)$ The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the number of logical operators $k$ for a fixed $nJ=240$. $b)$ Ratio between DM and QNDM resources numbers $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
The results of Fig. 5 illustrate the expected advantage of QNDM over the DM approach. Panel a), on the left, shows the resource cost $\mathcal{C}\left(g_{j}^{i}\right)$ for the two approaches in the limit $k\gg nJ$. They are plotted as a function of the number of logical operators $k$ needed to implement $U(\vec{\theta})$. The simulations have been done for $n=10$ qubits and a with $J=24$ different Pauli strings. Each value in Fig. 5 is calculated by averaging over different realizations with randomly chosen values of the parameters $J$, $k$, and the set of Pauli strings in $\hat{M}$. The average is performed also over the unitary transformations $U(\vec{\theta})$. This is obtained through a series of layers parameterized by $U_{j}(\vec{\theta}_{j})$ and entanglement layers denoted as $V_{j}$, as illustrated in Eqs. (3) and (4). Within the parameterized layer, the constituent single-qubit rotations $R^{j}_{i}(\theta_{i}^{j})$, and their corresponding parameters $\vec{\theta}_{i}^{j}$, are selected at random.
In panel a) of Fig. 5, as expected, QNDM needs fewer resources than DM  to evaluate the derivatives. Less obvious is that already for a limited number of logical operators, i.e. $k\sim 10^{3}$, we save more than $10^{2}$ logical gates.
In panel b) of Fig. 5, on the right, we show the ratio of the resources in terms of the number of Pauli strings $J$, comparing different possible choices for $k$. Already for $k\sim 10^{3}$, the ratio approaches the theoretical prediction in Eq. (29) (shown as a dashed black curve in the plot) and the DM method needs more than $100$ times more resources than the QNDM one. Due to the increasing number of shots needed for the DM when $J$ increases, we expect it to reach already $10^{6}$ for simple practical applications Wecker et al. (2015).
Figure 6: Regime $k\ll nJ$. The averaged number of resources $\mathcal{C}(g_{j}^{i})$ needed to estimate the derivative of the cost function along the direction $j$ ($i=QNDM$(orange line) and $i=DM$(blue line)). $\mathcal{C}(g_{j}^{i})$ is plotted as a function of the product $nJ$ for a fixed number of logical operators $k=0.5\times 10^{3}$. $b)$ Ratio between DM and QNDM resources $\left(\mathcal{C}\left(g_{j}^{DM}/g_{j}^{QNDM}\right)\right)$ as a function of the number of logical operators $k$ each colored line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=500$, for DM, it is calculated for each realization by (28), such that the MSE errors are equal for both methods.
The regime in which $k\ll nJ$ is analyzed in Fig. 6. Panel a), on the left, shows the resource cost function $\mathcal{C}(g_{j}^{i})$ ($i=QNDM,DM$) obtained to estimate the derivative along the direction $j$. The values are plotted as a function $nJ$. The simulations are done for $n=10$ qubits with $k=0.5\cdot 10^{3}$ fixed. Each value in Fig. 6 is calculated by performing an average over a collection of different realizations, as discussed above. As delineated in tab. 1, for DM (the blue line) the number of resources increases quadratically in the number of $J$. Already for $kJ\sim 10^{3}$, the DM method need $10^{4}$ more resources than QNDM one.
In panel b), on the right, we show the ratio of the resources as a function of the number of logical operators $k$ and for different values of $nJ$. The expected advantage of the QNDM  approach is confirmed in this regime. Already for $nJ\sim 10^{3}$ the DM approach needs almost $10^{5}$ times more resources than the QNDM one.
Notice that the results obtained complement the analytic analysis provided in Solinas et al. (2023). Specifically, the error analysis, Sec. III and Sec. IV.1, shows that, for the QNDM, the statistical uncertainty, i.e., the mean square error, is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM, it increases linearly with $J$. This ensures additional savings in resources, as shown in Fig. 5 and Fig. 6, compared to what was anticipated in our analytical analysis Solinas et al. (2023). Additionally, in all our simulations and both regimes, we have considered relatively small and simple systems to provide a more conservative estimate. We might expect a further increasing advantage in realistic computational problems, as they usually require more logical gates and involve more complex observables $\hat{M}$.
V Second derivative and Hessian calculation
The QNDM approach can be naturally extended to the calculation of higher derivatives Solinas et al. (2023). Having access to the second derivatives of the cost function allows us to evaluate the Hessian matrix. This can be useful for the implementation of second-order optimizers like the Newton optimizer or the Diagonal Newton optimizer Mari et al. (2021).
With a straightforward extension of Eq. (7), it can be shown that the second derivative, requires the evaluation of the cost function $f$ in four points Mari et al. (2021); Solinas et al. (2023). To implement this in the DM approach, we need four independent measurements of the observables $\hat{M}$. On the other hand, following the QNDM paradigm, we can just couple the system to the detector four times to store in the phase of the latest the same amount of information Solinas et al. (2023). Then, analogously to the first-order derivative, we perform a measure on the detector only once. For further details and the numerical simulations about the second derivatives with QNDM method, we address the interested reader to Appendix V. In this section, instead, we present the persistence of the advantages observed in first derivatives even when extending to second derivatives.
V.1 Costs Analysis for second derivatives
In Table 2, we insert the costs for calculating second derivatives. We denote the resource cost function, $\mathcal{C}\left(g_{w,j}^{i}\right)$, for both the approaches ($i=QNDM,\,DM$) in terms of the number of gates required to compute the second derivative along the directions $w$ and $j$. These values are obtained from Refs. Solinas et al. (2023); Cerezo et al. (2021); Mari et al. (2021). As above, the costs depend on the number of Pauli strings $J$, the number of qubits $n$, the number of shots or repetitions $N_{i}$, and the number of logical operators needed to implement $U(\vec{\theta})$, denoted by $k$.
Method Shots Costs
DM $N_{DM}$ $\mathcal{C}\left(g_{w,j}^{DM}\right)=N_{DM}4J(k+n)$
QNDM $N_{QNDM}$ $\mathcal{C}\left(g_{w,j}^{QNDM}\right)=N_{QNDM}(7k+16Jn)$
Table 2: Resource cost function to compute second derivatives with the QNDM and the DM approaches.
From Table 2, we observe that the cost of calculating the second derivative with DM is twice the cost for the first derivative with the same method. For QNDM, the variation in cost is due to the coefficient of $k$ (which increases from $3$ to $7$) and to the coefficient of $J$, which is twice the one for a single derivative.
For a fixed value of MSE error and when $k\gg nJ$, the ratio of the resource cost functions reads
$\mathcal{C}\left(\frac{g_{w,j}^{DM}}{g_{w,j}^{QNDM}}\right)=\frac{4J}{7}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(J^{2})$ (31)
where for the second derivative for DM approach, to reach the same precision obtained with the QNDM approach, the number of shots is proportional to $N_{DM}\propto N_{QNDM}J$. This is shown in Appendix A. The opposite regime is $k\ll nJ$. In this case, the resource ratio is given by
$\mathcal{C}\left(\frac{g_{w,j}^{DM}}{g_{w,j}^{QNDM}}\right)=\frac{k}{4n}\frac{N_{DM}}{N_{QNDM}}=\mathcal{O}(kJ)$ (32)
These results demonstrate that the advantage in terms of the cost function for QNDM is also maintained in the case of the second derivative. As already mentioned, all the simulations used to validate these results are included in Appendix A.
VI Conclusions
In this paper, we have presented a detailed study of an alternative approach to evaluating the derivatives of a cost function using a quantum computer, along with its implementation in Python publicly available via GitHub S.Caletti and G.Minuto (2024). This method called Quantum Non-Demolition Measurement (QNDM), was first introduced in Ref. Solinas et al. (2023). In addition to providing a quantitative estimate of the resources needed to run the QNDM protocol, we have conducted an in-depth comparison with the state-of-the-art Direct Measurement (DM) approach Cerezo et al. (2021). The results obtained are encouraging and pave the way for the first applications of the QNDM  algorithm to real-world problems since it allows us to consistently reduce the resources needed for the computation.
This advantage can be attributed to two main features. First, as highlighted in Solinas et al. (2023), the use of a quantum detector to store information about the derivative of the cost function allows us to run the quantum circuit and measure only once. In contrast, the DM approach requires two circuits and two measurements (four circuits and four measurements for the Hessian). Second, introduced here for the first time, the statistical uncertainty (i.e., the mean square error) for the QNDM is independent of the number of Pauli strings $J$ appearing in the quantum observable, whereas for the DM approach, it increases linearly with $J$. Consequently, to achieve the same accuracy in the average values of the cost function for quantum observables composed of numerous Pauli strings $J$, the number of quantum circuit runs in the DM approach must increase.
The numerical simulations performed using the Qiskit framework have confirmed the analytical estimates. Even for small systems and simplified quantum circuits, we observe a considerable reduction in the resources needed to run the QNDM approach compared to the DM method. We anticipate that this advantage will further increase for more practical and complex problems, such as the simulation of chemical compounds, molecules, or small quantum systems Weber et al. (2014). These results position the QNDM approach as a valuable alternative for implementing Variational Quantum Algorithms (VQAs) on the next generation of noisy quantum computers.
Acknowledgements.
PS and GM acknowledge financial support from INFN. This work was carried out while GM was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with the Dept. of Informatics, Bioengineering, Robotics, and Systems Engineering, Polytechnic School of Genova.
SC would like to thank the University of Amsterdam and the Delta-ITP program for their support during the initial part of this work.
Appendixes
Appendix A Second derivatives
In this section, we present a detailed analysis for the second derivative. The parameter shift rule, Eq. (5), can be generalized to calculate the second derivative Mari et al. (2021).
$g_{w,j}$ $=$ $\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{j}\partial\theta_{w}}=$
$=$ $\Big{[}f(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))-f(\vec{\theta}+s(-\hat{e}_{j}+\hat{e}_{w}))$
-- $f(\vec{\theta}+s(\hat{e}_{j}-\hat{e}_{w}))+f(\vec{\theta}-s(\hat{e}_{j}+\hat{e}_{w}))\Big{]}\Big{[}2\sin^{2}s\Big{]}^{-1}.$
We can re-write Eq. (A) in terms the density matrix representing the $n$-qubits initial state, i.e. $\rho_{s}^{0}=|\psi_{0}\rangle\langle\psi_{0}|$
$\frac{\partial^{2}f({\vec{\theta}})}{\partial\theta_{j}\partial\theta_{i}}$ $=$ $\Big{[}{\rm Tr}_{S}[U^{\dagger}(\vec{\theta}+s(e_{j}+e_{w}))\hat{M}U(\vec{\theta}+s(e_{j}+e_{w}))\rho_{S}^{0}-U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))\hat{M}U(\vec{\theta}+s(-e_{j}+e_{w}))\rho_{S}^{0}$
-- $U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))\hat{M}U(\vec{\theta}+s(e_{j}-e_{w}))\rho_{S}^{0}+U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))\hat{M}U(\vec{\theta}-s(e_{j}+e_{w}))\rho_{S}^{0}]\Big{]}\Big{[}4\sin^{2}s\Big{]}^{-1}.$
It follows that for DM method to calculate the values of $f(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))$, we run the quantum circuit to implement the corresponding $U(\vec{\theta}+s(\hat{e}_{j}+\hat{e}_{w}))$ and then perform a projective measure for each Pauli string composing the observable $\hat{M}$ Mari et al. (2021). The same procedure is then implemented for $f(\vec{\theta}+s(-\hat{e}_{j}+\hat{e}_{w})),f(\vec{\theta}+s(\hat{e}_{j}-\hat{e}_{w})),f(\vec{\theta}-s(\hat{e}_{j}+\hat{e}_{w}))$, then the derivative can be easily extracted. In Fig. S1 the DMis represented schematically.
Figure S1: Quantum circuit of the implementation of the DM protocol. Here $U(\vec{\theta}+s\hat{e}_{j}+s^{\prime}\hat{e}_{w})$ is Eq. (1) with two parameters shifted in directions $\hat{e}_{j}$ and $\hat{e}_{w}$. The $J$ is equal to the number of Pauli strings and the values ±$\pm s$ and ±$\pm s^{\prime}$ are the shift for the parameter shift rule.
On the other hand, in the QNDM method Solinas et al. (2023), we estimate the second derivative measuring the circuit shown in Fig. S2, and we avoid the repetitions for each Pauli string and for each term of the parameter shift rule. The unitary transformation corresponding to the full QNDM evolution is
$U^{2}_{tot}=e^{i\lambda Z_{a}\otimes\hat{M}}U_{4}e^{-i\lambda Z_{a}\otimes\hat{M}}U_{3}e^{i\lambda Z_{a}\otimes\hat{M}}U_{2}e^{-i\lambda Z_{a}\otimes\hat{M}}U_{1},$ (S3)
where the $U_{y}$ operators (with $y=1,..,4$) correspond to the operators working in the $\vec{\theta}$ space and they read
$U(\vec{\theta}+s(e_{j}-e_{w}))$ $=$ $U_{1}$
$U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))U(\vec{\theta}-s(e_{j}+e_{w}))$ $=$ $U_{2}$
$U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))U(\vec{\theta}+s(-e_{j}+e_{w}))$ $=$ $U_{3}$
$U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))U(\vec{\theta}+s(e_{j}+e_{w}))$ $=$ $U_{4}.$ (S4)
Figure S2: Quantum circuit of the implementation of the QNDM protocol for second derivatives. Here, $H$ is the Hadamard gate, $S$ is the phase gate, $U_{1}=U(\vec{\theta}+s(e_{j}-e_{w}))$, $U_{2}=U^{\dagger}(\vec{\theta}+s(e_{j}-e_{w}))U(\vec{\theta}-s(e_{j}+e_{w}))$, $U_{3}=U^{\dagger}(\vec{\theta}-s(e_{j}+e_{w}))U(\vec{\theta}+s(-e_{j}+e_{w}))$, $U_{4}=U^{\dagger}(\vec{\theta}+s(-e_{j}+e_{w}))U(\vec{\theta}+s(e_{j}+e_{w}))$ and $U_{\pm}=\exp\{\pm i\lambda\hat{Z_{a}}\otimes\hat{M}\}$ is the system-detector coupling operator.
A.1 Errors Analysis for Second Derivative
The MSEs for the second derivatives for both methods, respectively, are equal to
$\text{MSE}_{QNDM}(\hat{g}_{w,j})$ $=\frac{\lambda^{2}(\partial^{2}_{\lambda}\mathcal{G}_{\lambda})^{2}}{4}+\frac{\sigma_{D}^{2}}{16N_{QNDM}\sin^{4}s}\frac{1}{\lambda^{2}(1-(2P_{0}-1)^{2})}$ (S5)
$\text{MSE}_{DM}(\hat{g}_{w,j})$ $=\sum_{i}\frac{h^{2}_{i}\sigma_{s}^{2}}{4N_{DM}\sin^{4}s}.$ (S6)
Notice that the differences with the first derivative MSEs in Eqs. (27) and (23) are only in the denominator and for the second derivative, the MSEDM grows linearly with the number of Pauli strings $J$. Consequently, to achieve the same precision level as MSEQNDM in the DM approach, we need to increase the number of shots as a function of $J$. More specifically, we have that
$\frac{\text{MSE}_{DM}(\hat{g}_{w,j})}{\text{MSE}_{QNDM}(\hat{g}_{w,j})}=1\longrightarrow N_{DM}=N_{QNDM}\frac{4\sigma_{s}^{2}\lambda^{2}(1-(2P_{0}-1)^{2})}{\sigma^{2}_{D}}\sum_{i}h^{2}_{i}.$ (S7)
Where, similarly to Eq. (28), we keep the linearity in $J$, only with a modified multiplicative factor due to the differences in the denominators between the MSE for the first and the second derivatives.
A.2 Cost Simulations for Second Derivative
In Section V.1, we found the asymptotic cost ratios $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ for the two regimes: $k\ll nJ$ Eq. (32) and $k\gg nJ$ Eq. (31). In Figs. S4 and S3, we include simulations that illustrate how the theoretical results are also maintained for the second derivatives. For each value presented in the figures, we averaged over $L=50$ different realizations.
Figure S3: Regime $k\gg nJ$. $a)$ Number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of logical operators $k$ for a fixed $nJ=300$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of Pauli string $J$ each colored line corresponds at a fixed $k$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S3 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in the function of the logical operators $k$, when we consider the limit of $k\gg nJ$. The simulations are done for $n=10$ qubits, fixed $nJ=300$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of the number of Pauli strings $J$ and for $k$ different numbers of logical operators in $U(\vec{\theta})$.
Figure S4: $a)$ The averaged number of resources $\mathcal{C}(g_{w,j}^{i})$ needed to estimate the second derivative of the cost function along the directions $j$ and $w$; here, $i=QNDM$ (orange line) and $i=DM$ (blue line). $\mathcal{C}(g_{w,j}^{i})$ is plotted in function of the number of Pauli strings $J$ for a fixed number of logical operators $k=0.5\cdot 10^{3}$. $b)$ Ratio between DM and QNDM resources numbers $\mathcal{C}\left(g_{w,j}^{DM}/g_{w,j}^{QNDM}\right)$ as a function of logical operators $k$ each coloured line corresponds at a fixed $nJ$. The quantum circuit is composed of $n=10$ qubits and the average is taken over $L=50$ realization as discussed in the main text. The number of shots for the QNDMis $N_{QNDM}=200$, for DM, it is calculated for each realization by (S7), such that the MSE errors are equal for both methods.
Panel a) in Fig. S4 shows the numbers of resources, $\mathcal{C}\left(g_{w,j}^{i}\right)$, in function of $nJ$. We are working in the limit $k\ll nJ$. The simulations are done for $n=10$ qubits, a fixed number of logical operators $k=0.5\cdot 10^{3}$, and fixed shots $N=200$. In the second panel b) we show the ratio of the resources as a function of logical operators $k$ for different values of the product $nJ$. The advantage of the QNDM over the DM approach is confirmed also for the second derivatives.
Appendix B how to install QNDM
To install the QNDM library just clone the GitHub repository (link in S.Caletti and G.Minuto (2024)) using the command
git clone https://github.com/simonecaletti/qndm-gradient.git
All the functions for the QNDM algorithm are defined in the qndm folder. The test_scripts folder instead contains a set of scripts with some simple tests, like QNDM and DM derivatives evaluation and optimization tasks. To have access to the interface with the hamlib library Sawaya et al. (2024) you need to install the mat2qubit package Sawaya (2022). The instructions are contained in the hdf5-install.sh script, so just run
bash hdf5-install.sh
HamLib (for Hamiltonian Library), is freely available online and contains problem sizes ranging from 2 to 1000 qubits. It includes problem instances of the Heisenberg model, Fermi-Hubbard model, Bose-Hubbard model, molecular electronic structure, molecular vibrational structure, MaxCut, Max-k-SAT, Max-k-Cut, QMaxCut, and the traveling salesperson problem Sawaya et al. (2024). To test the installation create the folder output_test and run a test script, for example
python3 test_qndm.py
Run it from the qndm-gradient/ folder or add the corresponding path to your PYTHON PATH environment variable. If the installation is working correctly a QNDM_der.csv and a RunCard_Der.txt file have been created. The first one contains information about the gradient computation using the QNDM algorithm, while the second is an automatically generated runcard containing the details of the run.